OpenStudy (amtran_bus):

Calculate the wavelength of light, in nm, illuminating a barium surface if the threshold frequency of barium is 6.07x10^14 HZ and the kinetic energy of the electron emitted is 1.31x10^-18.

OpenStudy (amtran_bus):

@timo86m @karatechopper

OpenStudy (anonymous):

\[\phi _{Barium}=E _{light}-E _{electron}\] \[\phi _{Barium}=hf _{Barium}\] \[E _{light}=hf _{light}\]

OpenStudy (anonymous):

\[hf _{Ba}=hf _{Light}-E _{electron} \]

OpenStudy (anonymous):

\[c=\lambda f \rightarrow f=c/\lambda\] \[hf _{Ba}=hc/\lambda_{light}-E _{electron}\] solve for lambda, the wavelength in meters: \[\lambda _{light}= \frac{ hc}{hf _{Ba}+E _{electron}} \] h = 6.63*10^-34 J*s c = 3.00*10^8 m/s fba = 6.07*10^14 1/s Eelectron=1.31x10^-18 J Plug in, solve, and convert to nm by multiplying by (10^9 nm / 1 m) Answer should be 116 nm

OpenStudy (amtran_bus):

You are a life saver!

OpenStudy (amtran_bus):

Thank you very much.

OpenStudy (anonymous):

np it was a good refresher